# On regions of mixed unitarity for semigroups of unital quantum channels

### v1

@@ 2.1. Notations
Since $\mathscr{MU}(d)$ is a closed convex set, we have $\mathcal{M}$ is a closed convex subset of a real vector space $\mathcal{N}$. Suppose $\Phi\notin\mathscr{MU}(d)$, then $\mathscr{J}(\Phi-\delta_{d})\in\mathscr{J}(\mathcal{N})\setminus\mathscr{J}(\mathcal{M})$. Note that any linear functional on $\mathscr{J}(\mathcal{N})$ is of the form $Z\mapsto\operatorname{\textsf{tr}}(YZ)$ for some $Y\in\mathscr{J}(\mathcal{N})$. Therefore, by the Hahn-Banach separation theorem  [^1] Theorem [Ze12, Theorem 1. C]: Let $M$ be a closed non-empty convex subset of a normed linear space $\mathcal{X}$ over $\mathbb{R}$. Let $u_{0}\in\mathcal{X}$ but not in $M$, then there exists a continuous linear function $f$ on $\mathcal{X}$ such that $f(u)\leq 1<f(u_{0})$ for all $u\in M$. there exist $Y\in\mathscr{J}(\mathcal{N})$ such that
@@
$(iv)$ Suppose the interior of $\mathscr{MU}(d)$ is non-empty. i.e., there exists a $\Phi_{0}\in\mathscr{MU}(d)$ and $r>0$ such that $B(\Phi_{0},r):=\{\Psi\in\mathscr{H}(d):\left\lVert\Psi-\Phi_{0}\right\rVert_{2}<r\}$ [^2] Since any two norms are equivalent in finite dimensional space, so we can choose the ball in Hilbert-Schmidt norm. is a subset of $\mathscr{MU}(d)$. Here $\left\lVert\Psi\right\rVert_{2}:=\operatorname{\textsf{tr}}\big(\mathscr{J}(\Psi)^{*}\mathscr{J}(\Psi)\big)$. Observe that $\mathscr{J}(B(\Phi_{0},r)-\delta_{d})\subseteq\mathcal{X}$, where $\mathcal{X}=\{\rho\in(\mathbb{M}_{d}\otimes\mathbb{M}_{d})_{Her}:\operatorname{\textsf{id}}\otimes\operatorname{\textsf{tr}}(\rho)=\operatorname{\textsf{tr}}\otimes\operatorname{\textsf{id}}(\rho)=0\}$ and $\delta_{d}=\frac{\operatorname{\textsf{tr}}(.)}{d}I_{d}$. Since $\mathscr{J}$ is an isometry and the translation of an open set is open, we deduce that $\mathcal{X}$ contains an open set. Therefore, $\mathcal{X}=(\mathbb{M}_{d}\otimes\mathbb{M}_{d})_{Her}$, which is not possible. ∎
@@ 3. Discrete semigroups of mixed unitary channels
Example 2.5 shows that for $d\geq 3$, not every unital quantum channel $\Phi$ on $\mathbb{M}_{d}$ is a mixed unitary. However, it is proven recently that the finite evolution of the unital quantum channel would lead to a mixed unitary at some point. More explicitly Theorem 21 of [KJRM23] shows that if $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is a unital quantum channel, then there exists some $k\in\mathbb{N}$ such that $\Phi^{k}$ is a mixed unitary channel. This suggests the following definition.
@@
Clearly, if $\Phi^{k}$ is a mixed unitary so is $\Phi^{mk}$ for all $m\geq 1.$ Hence the main result of [KJRM23] is telling us that for any unital quantum channel $\Phi$, $\{\Phi^{n}\}_{n\in\mathbb{N}}$ is frequently mixed unitary. In this article (see Theorem 3.9), we make the stronger assertion that it is eventually mixed unitary. This was known in some special cases, such as primitive (meaning that it is irreducible and has a trivial peripheral spectrum) channels ([HRS20, KJRM23]) and here we have it in general. The main tool we need is a result from [KJRM23] which says that quantum channels fixing elements of a subalgebra $\mathcal{A}$ and sufficiently close to the conditional expectation map on to $\mathcal{A}$ are mixed unitary. We begin with relevant definitions and results.
@@
Let $\mathcal{A}$ be a unital $C^{*}$-subalgebra of $\mathbb{M}_{d}$. A CP map $\mathbf{E}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ with $\mathbf{E}(\mathbb{M}_{d})=\mathcal{A}$ is said to be a conditional expectation onto $\mathcal{A}$ if $\mathbf{E}(A)=A,\forall A\in\mathcal{A}$.
Note that a conditional expectation map $\mathbf{E}$ as above is of norm 1 as it is a unital CP map and it is also a projection as it is the identity map on its range. It is also known that $\mathbf{E}(AXB)=A\mathbf{E}(X)B$ for all $A,B\in\mathcal{A}$ and $X\in\mathbb{M}_{d}.$ See [Ume53, Ra06].
Conditional expectation onto a $C^{*}$ subalgebra need not be unique. However, there is only one conditional expectation that preserves the trace, denote it by $\mathbf{E}_{\mathcal{A}}$. In fact, it is equal to the orthogonal projection of $\mathbb{M}_{d}$ onto $\mathcal{A}$ w.r.t. the Hilbert-Schmidt inner product on $\mathbb{M}_{d}$. In [BhBha25] one can see a concrete way of writing $\mathbf{E}_{\mathcal{A}}$ as a mixed unitary channel. One of the fundamental quantum channels widely discussed in the literature is the completely depolarizing channel $\delta_{d}$:
$\delta_{d}(X)=\frac{1}{d}\operatorname{\textsf{tr}}(X)I_{d},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d}.$ (3.1)
Note that this is the unique trace-preserving conditional expectation onto the algebra $\mathbb{C}I_{d}$. The following Proposition is due to John Watrous[Wat09]. It says that all unital quantum channels sufficiently close to the depolarizing channel are mixed unitary.
@@
[Wat09, Corollary 2] Fix $d\in\mathbb{N}.$ There exists $r>0$ depending on $d$ such that, every unital, quantum channel $\Phi$ on $\mathbb{M}_{d}$ with $\left\lVert\Phi-\delta_{d}\right\rVert<r$, is mixed unitary.
In 2024, David Kribs et al. [KJRM23] extended this result with the depolarising channel replaced by conditional expectation maps on arbitrary unital $C^{*}$-subalgebras of $\mathbb{M}_{d}$. For a given a unital $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$, we let $\mathcal{F}(\mathcal{A})$ to be the set of all unital quantum channels which leave elements of $\mathcal{A}$ fixed. That is,
@@
It is easy to see that $\mathcal{F}(\mathcal{A})$ is closed under the compositions. Furthermore, it is closed under adjoints also. This follows from the fact that the fixed point space of $\Phi$ is equal to the commutant of the set $\{V_{j}:1\leq j\leq\ell\}$, where $\{V_{j}\}_{j=1}^{\ell}$ is any set of Kraus operators of $\Phi$ (See [Kri03, Theorem 2.1]).
The following result has been observed and used in [KJRM23]. However, it is not explicitly stated or proved. The result is essential for us and so we provide the details.
@@
[KJRM23, Theorem 4, Lemma 12] Let $\mathcal{A}$ be a unital $C^{*}$-subalgebra of $\mathbb{M}_{d}$ and let $\mathbf{E}_{\mathcal{A}}$ be the unique trace-preserving conditional expectation onto $\mathcal{A}$. Then there exists $r_{\mathcal{A}}>0$ such that, if $\Phi\in\mathcal{F}(\mathcal{A})$ with $\left\lVert\Phi-\mathbf{E}_{\mathcal{A}}\right\rVert\leq r_{\mathcal{A}}$, then $\Phi$ is mixed unitary.
@@
Lemma 12 of [KJRM23] proves that $\mathbf{E}_{\mathcal{A}}$ is mixed unitary. Theorem 4 of the same article shows that there exists a $p\in(0,1)$ such that $p\Phi+(1-p)\mathbf{E}_{\mathcal{A}}$ is mixed unitary, for all $\Phi\in\mathcal{F}(\mathcal{A})$. Now, for a given $\Phi\in\mathcal{F}(\mathcal{A})$, define $\Phi_{p}:=\mathbf{E}_{\mathcal{A}}+\frac{1}{p}(\Phi-\mathbf{E}_{\mathcal{A}})$. Clearly, $\Phi_{p}$ fixes elements of $\mathcal{A}$. Suppose $\Phi_{p}$ is CP, then $\Phi_{p}\in\mathcal{F}(\mathcal{A})$, and hence $p\Phi_{p}+(1-p)\mathbf{E}_{\mathcal{A}}=\Phi$ is mixed unitary. Therefore, it suffices to prove that there exists $r_{\mathcal{A}}\geq 0$ such that $\left\lVert\Phi-\mathbf{E}_{\mathcal{A}}\right\rVert\leq r_{\mathcal{A}}$, then $\Phi_{p}$ is CP. For this, without loss of generality assume that $\mathcal{A}=\bigoplus_{k=1}^{\ell}(I_{m_{k}}\otimes\mathbb{M}_{n_{k}})$, where
@@
and the Choi-matrix of $\Phi$ is given as
@@
$e_{rr^{\prime}}^{(j,k)},f_{ss^{\prime}}^{(j,k)}$ are the standard matrix units of $\mathbb{M}_{m_{j}\times m_{k}},\mathbb{M}_{n_{j}\times n_{k}}$, respectively. Note that if $\mathcal{K}=\mathcal{K}_{1}\oplus\mathcal{K}_{2}$, then the matrix units $e_{rr^{\prime}}$ in $\mathscr{B}({\mathcal{K}_{1}})$ is treated as $E_{rr^{\prime}}=\begin{bmatrix}e_{rr^{\prime}}&0\\ 0&0\end{bmatrix}$ in $\mathscr{B}({\mathcal{K}})$. Now, by swapping the second and third tensor products in each block entry, the Choi-matrix is equal to
@@
Thus, by letting $\Psi^{\prime}:=(m_{1}+m_{2})\Psi$, the Choi-matrix (upto permutation equivalence) of $\Phi$ is equal to the tensorial Schur product (see [VsSu15]) of $\frac{1}{d}\mathscr{J}(\Psi^{\prime})$ and $|{f}\rangle\langle{f}|$. That is,
@@
Therefore, the Choi-matrix (upto permutation equivalence) of $\Phi_{p}$ is $\frac{1}{d}\big(Z+\frac{1}{p}(\mathscr{J}(\Psi^{\prime})-Z)\big)\circ^{\otimes}|{f}\rangle\langle{f}|,$ where $Z=\begin{bmatrix}\frac{I_{m_{1}}\otimes I_{m_{1}}}{m_{1}}&0\\ 0&\frac{I_{m_{2}}\otimes I_{m_{2}}}{m_{2}}\end{bmatrix}$. Now, as $Z$ is strictly positive, taking $r$ as the smallest positive eigenvalue of $Z$, $Z+\Delta$ is positive, whenever $\Delta=\Delta^{*}$ with $\left\lVert\Delta\right\rVert\leq r$. Since considering a principal submatrix can only reduce the norm, we get $\left\lVert\mathscr{J}(\Phi)-\mathscr{J}(\mathbf{E}_{\mathcal{A}})\right\rVert\geq\frac{1}{d}\left\lVert\mathscr{J}(\Psi^{\prime})-Z\right\rVert$. Set $r_{\mathcal{A}}=\frac{pr}{d}$, then whenever $\Phi\in\mathcal{F}(\mathcal{A})$ with $\left\lVert\mathscr{J}(\Phi)-\mathscr{J}(\mathbf{E}_{\mathcal{A}})\right\rVert\leq r_{\mathcal{A}}$ we have
@@
Hence $\big(Z+\frac{1}{p}(\mathscr{J}(\Psi^{\prime})-Z)\big)\geq 0$. By [VsSu15, Proposition 1.3], $\frac{1}{d}\big(Z+\frac{1}{p}(\mathscr{J}(\Psi^{\prime})-Z)\big)\circ^{\otimes}|{f}\rangle\langle{f}|$ is positive. Further, this is equivalent to the map $\Phi_{p}$ being CP. This completes the proof. ∎
Definition 3.5.
A linear map $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ said to be a conditional *-automorphism if $\Phi=\mathbf{E}_{\mathcal{A}}\circ\Pi$, where $\mathbf{E}_{\mathcal{A}}$ is a trace-preserving conditional expectation onto some $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ and $\Pi$ is a *-automorphism on $\mathbb{M}_{d}$ which commutes with $\mathbf{E}_{\mathcal{A}}$.
In general an automorphsim of a subalgebra $\mathcal{A}$ of $M_{d}$ need not be implemented by a unitary. For instance, take $\mathcal{A}=\{a(E_{1}1+E_{22})+bE_{33}:a,b\in\mathbb{C}\}\subseteq\mathbb{M}_{3}$, where $E_{ij}$ denotes matrix units and consider the automorphism $a(E_{11}+E_{22})+bE_{33}\mapsto b(E_{11}+E_{22})+aE_{33}.$ However, there are no such issues if the automorphism preserves the trace as the following theorem shows.
Theorem 3.6.
Let $\mathcal{A}$ be a unital $C^{*}$-subalgebra of $\mathbb{M}_{d}$. Let $\Psi$ be a trace-preserving *-automorphism on $\mathcal{A}$, then there exists a unitary $U\in\mathbb{M}_{d}$ such that $\Psi=\mathrm{Ad}_{U}$.
@@
It is to be noted that generally the unitary $U$ of this theorem is not an element of the algebra $\mathcal{A}.$ However, $\mathrm{Ad}_{U}$ leaves $\mathcal{A}$ and $\mathcal{A}^{\perp}$ invariant. There is extensive literature on the asymptotics of CP maps and quantum channels. We will follow mostly [BhSa23]. We cite [AFK25] for recent work and the citations in it for further information.
@@
    $\mathscr{M}_{\Phi}:=\{X\in\mathbb{M}_{d}:\Phi(XY)=\Phi(X)\Phi(Y)\mbox{ and }\Phi(YX)=\Phi(Y)\Phi(X)\quad\penalty 10000\ \forall Y\in\mathbb{M}_{d}\}.$
@@
It is known that for a unital CP map $\Phi$, we have
Recall that for any CP map $\Phi$ on $\mathbb{M}_{d}$, we have $\Phi(X^{*})\Phi(X)\leq\left\lVert\Phi(I)\right\rVert\Phi(X^{*}X)$, for all $X\in\mathbb{M}_{d}$ (This inequality is known as the Kadison-Schwarz inequality). As a consequence of this inequality, we get that if $\Phi$ is a unital quantum channel, then the peripheral space $\mathcal{P}(\Phi)\subseteq\mathscr{M}_{\Phi}$.
@@
Let $\Phi$ be a unital quantum channel on $\mathbb{M}_{d}.$ Then $\{\Phi^{n}\}_{n\in\mathbb{N}}$ is asymptotically a conditional *-automorphism. i.e., there exists a conditional *-automorphishm $\alpha$ on $\mathbb{M}_{d}$ such that
@@
Also there exists an automorphism $\gamma$ of $\mathbb{M}_{d}$, such that $(\gamma\circ\Phi)^{n}$ converges to the conditional expectation map onto the peripheral space of $\Phi$ as $n$ tends to $\infty.$
@@
Let $\mathcal{P}(\Phi)$ be the peripheral space of $\Phi$. That is, $\mathcal{P}(\Phi)=\operatorname{\textsf{span}}\{X:\Phi(X)=\lambda X,\left\lvert\lambda\right\rvert=1\}$. Then, the assumption that $\Phi$ is trace-preserving ensures that [BhSa23], $\mathcal{P}(\Phi)$ is a $C^{*}$-sub algebra of $\mathbb{M}_{d}.$ Let $\mathbf{E}_{\mathcal{P}}$ be the trace-preserving conditional expectation onto $\mathcal{P}(\Phi)$. Considering $\mathbb{M}_{d}$ as a Hilbert space with Hilbert-Schmidt inner product, $\mathbf{E}_{\mathcal{P}}$ is just the orthogonal projection on to the subspace $\mathcal{P}(\Phi).$ Moreover we have the following ([BhSa23], Theorem 3.1 and 3.4):
@@
    $\mathbb{M}_{d}=\mathcal{P}(\Phi)\oplus\mathcal{N}(\Phi)$, where $\mathcal{N}(\Phi)=\{Y:\lim_{n\to\infty}\Phi^{n}(Y)=0\}$ is an orthogonal direct sum with respect to the Hilbert-Schmidt inner product;
@@
    $\Phi=\alpha+\beta$, where $\alpha:=\Phi\circ\mathbf{E}_{\mathcal{P}}:\mathbb{M}_{d}\to\mathcal{P}(\Phi)\subseteq\mathbb{M}_{d}\ \text{ and }\ \beta:=\Phi-\alpha=\Phi\circ(\operatorname{\textsf{id}}-\mathbf{E}_{\mathcal{P}}):\mathbb{M}_{d}\to\mathcal{N}(\Phi)\subseteq\mathbb{M}_{d}$;
@@
[…]
This shows that the map, $\alpha=\mathrm{Ad}_{U}\circ\mathbf{E}_{\mathcal{P}}=\mathbf{E}_{\mathcal{P}}\circ\mathrm{Ad}_{U}$ is a conditional *-automorphism. Furthermore,

### v2

@@ 2.1. Notations
Since $\mathscr{MU}(d)$ is a closed convex set, we have $\mathcal{M}$ is a closed convex subset of a real vector space $\mathcal{N}$. Suppose $\Phi\notin\mathscr{MU}(d)$, then $\mathscr{J}(\Phi-\delta_{d})\in\mathscr{J}(\mathcal{N})\setminus\mathscr{J}(\mathcal{M})$. Note that any linear functional on $\mathscr{J}(\mathcal{N})$ is of the form $Z\mapsto\operatorname{\textsf{tr}}(YZ)$ for some $Y\in\mathscr{J}(\mathcal{N})$. Therefore, by the Hahn-Banach separation theorem there exist $Y\in\mathscr{J}(\mathcal{N})$ such that
@@
$(iv)$ Suppose the interior of $\mathscr{MU}(d)$ is non-empty. i.e., there exists a $\Phi_{0}\in\mathscr{MU}(d)$ and $r>0$ such that $B(\Phi_{0},r):=\{\Psi\in\mathscr{H}(d):\left\lVert\Psi-\Phi_{0}\right\rVert_{2}<r\}$ [^1] Since any two norms are equivalent in finite dimensional space, so we can choose the ball in Hilbert-Schmidt norm. is a subset of $\mathscr{MU}(d)$. Here $\left\lVert\Psi\right\rVert_{2}:=\operatorname{\textsf{tr}}\big(\mathscr{J}(\Psi)^{*}\mathscr{J}(\Psi)\big)$. Observe that $\mathscr{J}(B(\Phi_{0},r)-\delta_{d})\subseteq\mathcal{X}$, where $\mathcal{X}=\{\rho\in(\mathbb{M}_{d}\otimes\mathbb{M}_{d})_{Her}:\operatorname{\textsf{id}}\otimes\operatorname{\textsf{tr}}(\rho)=\operatorname{\textsf{tr}}\otimes\operatorname{\textsf{id}}(\rho)=0\}$ and $\delta_{d}$ be the depolarizing channel as in (1.2). Since $\mathscr{J}$ is an isometry and the translation of an open set is open, we deduce that $\mathcal{X}$ contains an open set. Therefore, $\mathcal{X}=(\mathbb{M}_{d}\otimes\mathbb{M}_{d})_{Her}$, which is not possible. ∎
@@ 3. Discrete semigroups of mixed unitary channels
Example 2.4 shows that for $d\geq 3$, not every unital quantum channel $\Phi$ on $\mathbb{M}_{d}$ is a mixed unitary. However, Theorem 21 of [23] shows that if $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is a unital quantum channel, then there exists some $k\in\mathbb{N}$ such that $\Phi^{k}$ is a mixed unitary channel. This suggests the following definition.
@@
Clearly, if $\Phi^{k}$ is a mixed unitary so is $\Phi^{mk}$ for all $m\geq 1.$ Hence the main result of [23] is telling us that for any unital quantum channel $\Phi$, $\{\Phi^{n}\}_{n\in\mathbb{N}}$ is frequently mixed unitary. In this article (See Theorem 3.9), we make the stronger assertion that it is eventually mixed unitary. This was known in some special cases, such as primitive (meaning that it is irreducible and has a trivial peripheral spectrum) channels ([17, 23]) and here we have it in general. The main tool we need is a result from [23] which says that quantum channels fixing elements of a subalgebra $\mathcal{A}$ and sufficiently close to the conditional expectation map on to $\mathcal{A}$ are mixed unitary. We begin with relevant definitions and results.
@@
Let $\mathcal{A}$ be a unital $C^{*}$-subalgebra of $\mathbb{M}_{d}$. A CP map $\mathbf{E}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ with $\mathbf{E}(\mathbb{M}_{d})=\mathcal{A}$ is said to be a conditional expectation onto $\mathcal{A}$, if $\mathbf{E}(A)=A,\forall A\in\mathcal{A}$.
Note that a conditional expectation map $\mathbf{E}$ as above is of norm 1 as it is a unital CP map and it is also a projection ($\mathbf{E}^{2}=\mathbf{E}$) as it is the identity map on its range. It is also known that $\mathbf{E}(AXB)=A\mathbf{E}(X)B$ for all $A,B\in\mathcal{A}$ and $X\in\mathbb{M}_{d}.$ See [37, 32].
While a conditional expectation onto a $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ is not necessarily unique, there exists a unique trace-preserving conditional expectation. Denote it by $\mathbf{E}_{\mathcal{A}}$. It is equal to the orthogonal projection of $\mathbb{M}_{d}$ onto $\mathcal{A}$ with respect to the Hilbert-Schmidt inner product on $\mathbb{M}_{d}$. In [4] one can see a concrete way of writing $\mathbf{E}_{\mathcal{A}}$ as a mixed unitary channel. One of the fundamental quantum channels widely discussed in the literature is the completely depolarizing channel $\delta_{d}$ as defined in (1.2). Note that $\delta_{d}$ is the unique trace-preserving conditional expectation onto the algebra $\mathbb{C}I_{d}$. The following Proposition is due to John Watrous [38]. It says that all unital quantum channels sufficiently close to the depolarizing channel are mixed unitary.
@@
[38, Corollary 2] Fix $d\in\mathbb{N}.$ There exists $r>0$ depending on $d$ such that, every unital, quantum channel $\Phi$ on $\mathbb{M}_{d}$ with $\left\lVert\Phi-\delta_{d}\right\rVert<r$, is mixed unitary.
In 2024, David Kribs et al. [23] found a clever way of extending this result with the depolarising channel replaced by conditional expectation maps on to arbitrary unital $C^{*}$-subalgebras of $\mathbb{M}_{d}$. For a given a unital $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$, we let $\mathcal{F}(\mathcal{A})$ to be the set of all unital quantum channels which leave elements of $\mathcal{A}$ fixed. That is,
@@
It is easy to see that $\mathcal{F}(\mathcal{A})$ is closed under the compositions. Furthermore, it is closed under adjoints also. This follows from the fact that the fixed point space of $\Phi$ is equal to the commutant of the set $\{V_{j}:1\leq j\leq\ell\}$, where $\{V_{j}\}_{j=1}^{\ell}$ is any set of Kraus operators of $\Phi$ (See [24, Theorem 2.1]).
The following result has been observed and used in [23]. However, it is not explicitly stated or proved. The result is essential for us and so we provide the details.
@@
[23, Theorem 4, Lemma 12] Let $\mathcal{A}$ be a unital $C^{*}$-subalgebra of $\mathbb{M}_{d}$ and let $\mathbf{E}_{\mathcal{A}}$ be the unique trace-preserving conditional expectation onto $\mathcal{A}$. Then there exists $r_{\mathcal{A}}>0$ such that, if $\Phi\in\mathcal{F}(\mathcal{A})$ with $\left\lVert\Phi-\mathbf{E}_{\mathcal{A}}\right\rVert\leq r_{\mathcal{A}}$, then $\Phi$ is mixed unitary.
@@
Lemma 12 of [23] proves that $\mathbf{E}_{\mathcal{A}}$ is mixed unitary. Theorem 4 of the same article shows that there exists a $p\in(0,1)$ such that $p\Phi+(1-p)\mathbf{E}_{\mathcal{A}}$ is mixed unitary, for all $\Phi\in\mathcal{F}(\mathcal{A})$. Now, for a given $\Phi\in\mathcal{F}(\mathcal{A})$, define $\Phi_{p}:=\mathbf{E}_{\mathcal{A}}+\frac{1}{p}(\Phi-\mathbf{E}_{\mathcal{A}})$. Clearly, $\Phi_{p}$ fixes elements of $\mathcal{A}$. Suppose $\Phi_{p}$ is CP, then $\Phi_{p}\in\mathcal{F}(\mathcal{A})$, and hence $p\Phi_{p}+(1-p)\mathbf{E}_{\mathcal{A}}=\Phi$ is mixed unitary. Therefore, it suffices to prove that there exists $r_{\mathcal{A}}>0$ such that $\left\lVert\Phi-\mathbf{E}_{\mathcal{A}}\right\rVert\leq r_{\mathcal{A}}$, then $\Phi_{p}$ is CP. For this, without loss of generality assume that $\mathcal{A}=\bigoplus_{k=1}^{\ell}(I_{m_{k}}\otimes\mathbb{M}_{n_{k}})$, where
@@
The Choi-matrix of $\Phi$ is seen to be
@@
where $e_{rr^{\prime}}^{(j,k)},f_{ss^{\prime}}^{(j,k)}$ are the standard matrix units of $\mathbb{M}_{m_{j}\times m_{k}}$ and $\mathbb{M}_{n_{j}\times n_{k}}$ respectively. Note that if $\mathcal{K}=\mathcal{K}_{1}\oplus\mathcal{K}_{2}$, then the matrix units $e_{rr^{\prime}}$ in $\mathscr{B}({\mathcal{K}_{1}})$ is treated as $E_{rr^{\prime}}=\begin{bmatrix}e_{rr^{\prime}}&0\\ 0&0\end{bmatrix}$ in $\mathscr{B}({\mathcal{K}})$. Now, by swapping the second and third tensor products in each block entry, the Choi-matrix is equal to
@@
Thus, by letting $\Psi^{\prime}:=(m_{1}+m_{2})\Psi$, the Choi-matrix (up to permutation equivalence) of $\Phi$ is equal to the tensorial Schur product (See [35]) of $\frac{1}{d}\mathscr{J}(\Psi^{\prime})$ and $|{f}\rangle\langle{f}|$. That is,
@@
Therefore, the Choi-matrix (upto permutation equivalence) of $\Phi_{p}$ is $\frac{1}{d}\big(Z+\frac{1}{p}(\mathscr{J}(\Psi^{\prime})-Z)\big)\circ^{\otimes}|{f}\rangle\langle{f}|,$ where $Z=\begin{bmatrix}\frac{I_{m_{1}}\otimes I_{m_{1}}}{m_{1}}&0\\ 0&\frac{I_{m_{2}}\otimes I_{m_{2}}}{m_{2}}\end{bmatrix}$. Now, as $Z$ is strictly positive, taking $r$ as the smallest positive eigenvalue of $Z$, $Z+\Delta$ is positive, whenever $\Delta=\Delta^{*}$ with $\left\lVert\Delta\right\rVert\leq r$. Since a principal submatrix has norm less than or equal to the original matrix, we get $\left\lVert\mathscr{J}(\Phi)-\mathscr{J}(\mathbf{E}_{\mathcal{A}})\right\rVert\geq\frac{1}{d}\left\lVert\mathscr{J}(\Psi^{\prime})-Z\right\rVert$. Set $r_{\mathcal{A}}=\frac{pr}{d}$, then whenever $\Phi\in\mathcal{F}(\mathcal{A})$ with $\left\lVert\mathscr{J}(\Phi)-\mathscr{J}(\mathbf{E}_{\mathcal{A}})\right\rVert\leq r_{\mathcal{A}}$ we have
@@
Hence $\big(Z+\frac{1}{p}(\mathscr{J}(\Psi^{\prime})-Z)\big)\geq 0$. By [35, Proposition 1.3], $\frac{1}{d}\big(Z+\frac{1}{p}(\mathscr{J}(\Psi^{\prime})-Z)\big)\circ^{\otimes}|{f}\rangle\langle{f}|$ is positive. Further, this is equivalent to the map $\Phi_{p}$ being CP. This completes the proof. ∎
In general an automorphsim of a subalgebra $\mathcal{A}$ of $M_{d}$ need not be implemented by a unitary. For instance, take $\mathcal{A}=\{a(E_{11}+E_{22})+bE_{33}:a,b\in\mathbb{C}\}\subseteq\mathbb{M}_{3}$, where $E_{ij}$ denotes matrix units and consider the automorphism $a(E_{11}+E_{22})+bE_{33}\mapsto b(E_{11}+E_{22})+aE_{33}.$ However, there are no such issues if the automorphism preserves the trace as the following result shows.
Lemma 3.5.
Let $\mathcal{A}$ be a unital $C^{*}$-subalgebra of $\mathbb{M}_{d}$. Let $\Psi$ be a trace-preserving *-automorphism of $\mathcal{A}$. Then there exists a unitary $U\in\mathbb{M}_{d}$ such that $\Psi(X)=\mathrm{Ad}_{U}(X),\penalty 10000\ \penalty 10000\ \forall X\in\mathcal{A}$ and $\mathrm{Ad}_{U}$ leaves $\mathcal{A}^{\perp}$ invariant, where the orthogonal complement is taken with respect to the Hilbert-Schmidt inner product.
@@
Now we recall some basic notions required to study asymptotics of CP maps and quantum channels. Due to its importance in quantum theory of open systems, there is extensive literature on this topic. We will follow mostly [7]. However we refer the reader to [9, 21] for some recent works and the citations in them for further information.
@@
    $\mathscr{M}_{\Phi}:=\{X\in\mathbb{M}_{d}:\Phi(XY)=\Phi(X)\Phi(Y)\mbox{ and }\Phi(YX)=\Phi(Y)\Phi(X),\quad\penalty 10000\ \forall Y\in\mathbb{M}_{d}\}.$
@@
If $\Phi$ is a unital quantum channel, making use of Kadison-Schwarz inequality it can be seen that the peripheral space is contained in the multiplicative domain: $\mathcal{P}(\Phi)\subseteq\mathscr{M}_{\Phi}$.
Definition 3.7.
A linear map $\alpha:\mathbb{M}_{d}\to\mathbb{M}_{d}$ said to be a conditional *-automorphism if $\alpha=\Pi\circ\mathbf{E}_{\mathcal{A}}$, where $\mathbf{E}_{\mathcal{A}}$ is a trace-preserving conditional expectation onto some $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ and $\Pi$ is a *-automorphism on $\mathbb{M}_{d}$, which commutes with $\mathbf{E}_{\mathcal{A}}$.
@@
Let $\Phi$ be a unital quantum channel on $\mathbb{M}_{d}.$ Then $\{\Phi^{n}\}_{n\in\mathbb{N}}$ is asymptotically a conditional *-automorphism. i.e., there exists a conditional *-automorphism $\alpha$ on $\mathbb{M}_{d}$ such that
@@
Let $\mathcal{P}$ be the peripheral space of $\Phi$. That is, $\mathcal{P}=\operatorname{\textsf{span}}\{X:\Phi(X)=\lambda X,\left\lvert\lambda\right\rvert=1\}$. Then, the assumption that $\Phi$ is trace-preserving ensures that $\mathcal{P}$ is a $C^{*}$-sub algebra of $\mathbb{M}_{d}$ (See [7]). Let $\mathbf{E}_{\mathcal{P}}$ be the trace-preserving conditional expectation onto $\mathcal{P}$. Considering $\mathbb{M}_{d}$ as a Hilbert space with Hilbert-Schmidt inner product, $\mathbf{E}_{\mathcal{P}}$ is just the orthogonal projection on to the subspace $\mathcal{P}.$ Moreover we have the following ([7], Theorem 3.1 and 3.4):
@@
    $\mathbb{M}_{d}=\mathcal{P}\oplus\mathcal{N}$, where $\mathcal{N}=\{Y:\lim_{n\to\infty}\Phi^{n}(Y)=0\}$ is an orthogonal direct sum with respect to the Hilbert-Schmidt inner product;
@@
    $\Phi=\alpha+\beta$, where $\alpha:=\Phi\circ\mathbf{E}_{\mathcal{P}}:\mathbb{M}_{d}\to\mathcal{P}\subseteq\mathbb{M}_{d}\ \text{ and }\ \beta:=\Phi-\alpha=\Phi\circ(\operatorname{\textsf{id}}-\mathbf{E}_{\mathcal{P}}):\mathbb{M}_{d}\to\mathcal{N}\subseteq\mathbb{M}_{d}$;